- 4y = 71 +et+ sin(2) a. Find the general solution b. Find a solution such that y(0) = 0, 7(0) = 1. - 7
Answer #1

Solution :- dy - 44 7x tex + sin(x) dx2 las The y-49 = 7x+ ex + sin(x). auxilary equation is m2-4=0. - m² 2²0 => (M+2)(m-2)=P.I, - (0?=434 - ex .: P.1, - Yozycyzuet Replace o byl. pil, - Yaz lex). P: 5 : (D4) 3 : gio(*), :: P.1, = 1024 sioux) - Sin→ 4+- 13:01 ► G+C2 = 1/3 → (1), y(o)-1 -> 29 20-cyl° - 44 -% - Cosco) 2 l, => 24-222 = 1+74 +43 + Ys = 19460. => circa = 197